Sara left downtown Portland, and 3 hours later, sam left going 43mph faster to catch up. After another 2 hours, Sam caught up. Find saras average speed?
Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20060)

Let her average speed by r mph.

After 3 hours, since d=rt, she was r∙3 or 3r miles down the road 
when Sam left.

So the distance between them when Sam left was 3r.

His catch-up rate was 43 mph to catch up the 3r distance between them,
so since d=rt, 3r = 43∙2
               3r = 86
                r = 86/3 = 28 2/3 mph

Edwin

Answer by stanbon(75887)

Sara left downtown Portland. 3 hours later sam left going 43mph faster to catch up. After another 2 hours, Sam caught up. Find Sarah's average speed?
---
Sarah DATA:
rate = r mph ; time = 5 hrs ; distance = 5r miles
-----
Sam DATA:
rate = r+43 mph ; time = 2 hrs ; distance = (r+43)2 miles
--------------
Equation:
dist = dist
5r = 2r+86
---
3r = 86
r = 28 2/3 mph (Sarah's rate)
---
r + 43 = 71 2/3 mph (Sam's rate)
---------
Cheers,
Stan H.
-------------